OpenStudy (anonymous):

Solve for the variable in this equation. 1/4(2x-12)=2/3(3x+9)

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

But Y?

OpenStudy (anonymous):

k. what i did first was multiply the numbers inside the equations by the fractions, example: (1/4*2x)-(1/4*12)=(2/3*3x)+(2/3*9). This will give you: .5x-3=2x+6 Next, just get the variables on one side and the numbers on the other, example: .5x-.5x-3-6=2x-.5x+6-6 Combine like terms to get: -9=1.5x Now, simply divide both sides by 1.5 to get -6=x